OpenStudy (anonymous):

How to solve? cos( 2theta)+10 sin^2 theta =7

OpenStudy (zehanz):

I'll switch to x instead of theta, typing is a lot easier then... Your equation: cos2x+10sin²x=7. There is a (well-known) formula for cos2x: cos2x=1-2sin²x. If we put that in place of cos2x, we get: 1-2sin²x+10sin²x=7, which after cleaning up becomes: 8sin²x=6. I think it's your turn to do the next step now...